### v1

@@ 4.3 Influence of orbital parameters on tidal feature formation
To further investigate the orbital dynamical mechanisms that may drive the formation of tidal structures in globular clusters, we analyzed a set of orbital parameters that are commonly linked to tidal interactions with the Galactic potential. Specifically, we considered the pericentric distance ($R_{{\rm Peri}}$), apocentric distance ($R_{{\rm Apo}}$), orbital semi-major axis ($a$), orbital inclination angle ($i$), orbital eccentricity ($e$), and the radial angle in action-angle coordinates ($\theta_{\rm R}$). These quantities collectively describe the cluster’s trajectory through the Galaxy and its exposure to varying tidal forces over time. All basic data are taken from the globular cluster database888people.smp.uq.edu.au/HolgerBaumgardt/globular/orbits.html. The semi-major axis, eccentricity, and inclination are computed following the formulae in 2019ApJ...882...98P. The radial angle ($\theta_{\rm R}$) in action-angle coordinates is calculated using the Galpy package, adopting the MWPotential2014 Galactic potential (2015ApJS..216...29B). This parameter not only characterizes the current position of the cluster along its orbit but also indicates its direction of motion between the pericenter and apocenter. Specifically, when $0^{\circ}<\theta_{\rm R}<180^{\circ}$, the cluster is moving outward from pericenter toward apocenter, whereas when $180^{\circ}<\theta_{\rm R}<360^{\circ}$, it is moving inward from apocenter back toward pericenter.
@@ 4.3.1 Possible group-sensitive parameters
In the two-dimensional orbital parameter space, several notable correlations emerge. Clusters with smaller pericentric distances ($R_{\rm Peri}$) generally exhibit higher mass-loss fractions, consistent with the expectation that stronger tidal forces near the Galactic center accelerate mass depletion. This relation resembles a rough power-law trend, although some G3 clusters with small $R_{\rm Peri}$ show relatively low mass-loss fractions, suggesting that additional mechanisms may mitigate their long-term stripping. Clear geometric correlations are also evident among $R_{\rm Peri}$, $R_{\rm Apo}$, and $a$, but these simply reflect basic orbital mechanics. Clusters on highly eccentric ($e\gtrsim 0.7$) and strongly inclined ($i\gtrsim 70^{\circ}$) orbits are more frequently found among G1, and such orbital configurations are typically associated with enhanced mass loss, even though some overlap exists among different morphological groups. 2014MNRAS.445.1048W argued that repeated passages of globular clusters through the Galactic disk can amplify tidal disruption and increase mass loss. Building on this, 2019ApJ...882...98P proposed that clusters with higher eccentricities and inclinations are more likely to undergo repeated disk crossings, thereby increasing their susceptibility to tidal depletion. 2020A&A...637L...2P further showed that clusters on highly eccentric ($e\gtrsim 0.8$) and strongly inclined ($\approx\pm 70^{\circ}$) orbits tend to lose more mass than those on nearly circular, low-inclination orbits. Our results are broadly consistent with this framework, while also indicating that not all clusters on such orbits necessarily display prominent tidal features.
@@
Moreover, 2018MNRAS.474.2479B suggested that globular clusters exhibiting tidal tails are typically in a near-dissolution phase and are more likely to be found near apocenter. In our sample, a considerable fraction of G1 clusters are located in the outward-moving orbital phase between pericenter and apocenter (with $\theta_{\rm R}$ between 0 and 180), consistent with the expectation that enhanced tidal stripping occurs during pericentric passages. However, we do not find clear evidence that these clusters are preferentially concentrated near apocenter. Although $\theta_{\rm R}$ and the mass-loss fraction show no strong correlation, G1 clusters nonetheless tend to populate this outward-moving phase, suggesting that the tidal stripping experienced during recent pericentric passages may contribute to the formation of their observed extra-tidal structures.
@@
Taken together, our results indicate that certain orbital characteristics, such as smaller $R_{\rm Peri}$, higher $e$, and specific $\theta_{\rm R}$, may be associated with enhanced tidal mass loss, and that these three parameters are group-sensitive. These findings highlight that, beyond intrinsic properties, the orbital configuration of globular clusters also plays a key role in determining both the extent of their mass loss and the detectability of their tidal structures.
4.3.2 Tidal feature orientations vs. orbital direction
@@
Meanwhile, to investigate whether the orientation of tidal structures is related to the orbital direction of the clusters, we overlaid the orbital directions on the matched-filter S/N maps shown in Fig. 3. The sky-blue arrows indicate the short-period orbital directions derived from orbital integration, while the orange arrows mark the directions toward the Galactic center. The orbital integration was performed using the Galpy package with the MWPotential2014 gravitational model (2015ApJS..216...29B). This directional information provides additional context for interpreting whether the detected features are shaped primarily by orbital motion or Galactic tides.
@@
Our analysis reveals three distinct alignment patterns of tidal features among the clusters in our sample. In some systems (e.g., Whiting 1, NGC 5466), the extended structures broadly align with the orbital path, consistent with theoretical expectations that tidal stripping occurs preferentially along the orbit, producing leading and trailing tails (e.g., 2007ApJ...659.1212M; 2010MNRAS.401..105K). In others (e.g., NGC 1261, NGC 7492), the extensions are partially oriented toward the Galactic center, suggesting that the external tidal field has played a key role, possibly through recent disk-crossing events or similar perturbations (e.g., 1999A&A...352..149C; 2011MNRAS.416..393B). A third group of clusters (e.g., NGC 1851, NGC 6981) displays more symmetric, envelope-like morphologies with no clear alignment. In some cases, certain systems show visible tidal tails that are not aligned with the orbital direction (e.g., NGC 1904, NGC 4147), which may be related to recent pericentric passages (2025A&A...693A..69A).
@@
Figure 7: Corner plot of orbital parameters. The parameters included are: pericenter distance ($R_{\mathrm{Peri}}$), apocenter distance ($R_{\mathrm{Apo}}$), orbital semi-major axis ($a$), inclination angle(ii), orbital eccentricity($e$), and radial angle in action-angle coordinates ($\theta_{\rm R}$). The mass-loss fraction ($M_{\rm dis}/M_{\rm init}$) is also included to test its potential coupling with orbital properties. The color-coding of morphological classifications, the use of symbols, and the presentation of diagonal/off-diagonal panels are the same as in Fig. 6.
@@ 4.4 Environmental independence of tidal morphologies
Progenitor: To assess whether accretion origin plays a role in shaping the outer structures of globular clusters, we examined their distribution in the energy-angular momentum ($E$-$L_{z}$) plane (Fig. 9), a diagnostic that has been widely applied to trace the assembly history of the Galactic halo and to connect clusters with past merger events (see, e.g., 2019A&A...630L...4M; 2020ApJ...901...48N; 2022ApJ...926..107M). The associations between clusters and their progenitors adopted in this study are based on 2019A&A...630L...4M and 2020MNRAS.493..847F.
@@
Taken together, our analysis indicates that clusters with tidal features are distributed across a wide range of pericentric distances, orbital periods, pericentric passage times, and disk-crossing times, showing no clear preference for regions experiencing stronger or more frequent tidal forces. Likewise, accretion history provides no definitive evidence for driving the formation of extended tidal structures.
Figure 8: (a) The distribution of pericentric distance ($R_{\rm Peri}$) versus radial orbital period ($P_{\rm R}$). (b) The time since the last pericentric passage ($T_{\rm Peri}$) versus the time since the most recent disk crossing ($T_{\rm Disk}$). Clusters are color-coded by morphological group: G1 in red, G2 in blue, and G3 in black. Pal 12 is highlighted with an open marker.
Figure 9: Distribution of globular clusters in the $L_{z}$-$E$ space. Symbols indicate possible progenitor origins: circles represent clusters likely associated with the Gaia-Sausage-Enceladus (GSE) merger, triangles denote clusters linked to the Sagittarius (Sag) dwarf galaxy, squares indicate possible origin in the Sequoia (Seq) event, diamonds denote clusters associated with the Helmi stream (H99), and crosses mark clusters that may belong to either Gaia-Sausage-Enceladus or the Helmi stream (GSE/H99). Clusters are color-coded according to morphological classification: red for G1, blue for G2, and black for G3. Pal 12 is highlighted with open marker.
@@ 4.5 Limitations in the detection of tidal structures
Despite recent advances in wide-field surveys and data processing techniques, the identification of tidal structures around Galactic globular clusters remains subject to several observational and methodological limitations. These factors may introduce biases in the detection and interpretation of extended features, potentially leading to an underestimation of their prevalence, morphology, or physical origin.
  1. $\bullet$
@@
    The data depth is crucial for assessing the completeness of tidal structures. In this work, we make use of imaging data from the DESI Legacy Survey, which reaches down to $r$ or $g$ $\sim$23-24 mag. For the globular clusters in our sample, this depth corresponds to 2-3 magnitudes below the main-sequence turnoff, thereby encompassing the majority of stellar populations associated with the clusters. Importantly, the Legacy Survey represents the deepest publicly available wide-field dataset that contains the largest number of globular clusters. Thus, the data employed here greatly reduces the biases arising from limited photometric depth, allowing a more complete characterization of tidal features.
@@ 2. $\bullet$
    Methods such as matched filtering (2002AJ....124..349R) require careful choices of the spatial binning scale and the size of the smoothing kernel applied to density maps. Excessively large bins or overly broad kernels can make thin tidal tails difficult to discern, whereas overly small bins and kernels may introduce excessive noise and generate spurious structures. Therefore, by testing a range of spatial binning and smoothing scales and cross-checking our results against independent studies (see Appendix A), we minimize the risk of misclassification arising from parameter choices and ensure the reliability of our morphological classification.
  3. $\bullet$
@@
    The tidal boundary of a globular cluster is often approximated by the King model tidal radius(1962AJ.....67..471K), derived from fitting the surface density profile under the assumption of dynamical equilibrium. However, this estimate does not take into account orbital effects or the time-varying nature of the Galactic potential. Dynamical alternatives (such as the Jacobi radius and, in this work, the dynamical tidal radius at the current orbital phase) depend on the cluster’s mass, orbital phase, and the adopted Galactic potential model. Using different $r_{\rm t}$ values can change the classification of outer stars as ”bound” or ”extra-tidal”, thereby affecting whether a cluster is deemed to possess extended structures.
    To mitigate these effects, we adopted both the widely used $r_{\rm t}^{\rm King}$ (1996AJ....112.1487H 1996AJ....112.1487H, 2010 edition) from the literature and $r_{\rm t}^{\rm dyn}$ (from the globular cluster databases referenced in this work) calculated from orbital parameters. Both sets of values are taken from well-established and authoritative cluster catalogs, ensuring consistency with previous studies. In our morphological classification, we explicitly considered both definitions, thereby reducing the limitations associated with relying on a single tidal radius and enhancing the reliability of extra-tidal structure identification.
  4. $\bullet$
@@
In summary, although matched filtering and wide-field imaging have improved our ability to identify extended structures in globular clusters, the detection remains vulnerable to various systematic effects. In this study, we minimized these effects by employing a homogeneous dataset, adopting uniform filtering parameters and processing procedures, and cross-validating our results with previous imaging studies. These measures reduce potential systematic biases and enhance the reliability of the identified tidal features and their morphological classifications.
@@ 5 Summary
In this study, we conducted a systematic investigation of the extra-tidal structures associated with Galactic globular clusters of likely extragalactic origin, aiming to explore their formation mechanisms and the key physical factors driving their morphological diversity. Using a homogeneous dataset from the DESI Legacy Imaging Surveys and a uniform matched-filtering procedure, we identified and morphologically classified the tidal features of our cluster sample. We then explored the connections between extra-tidal structures and clusters’ intrinsic properties, orbital dynamical parameters, and Galactic environment, followed by a discussion of several observational and methodological limitations affecting their detection and classification. Within this framework, we arrived at the following results:
1. Our analysis shows that a substantial fraction of clusters exhibit extended tidal structures, including prominent tidal tails (G1) and diffuse stellar envelopes (G2), which in many cases extend well beyond both the King tidal radius and the dynamical tidal radius. In contrast, some clusters remain spatially confined within their nominal tidal boundaries and show no evidence of extra-tidal features (G3). In total, 12 clusters are classified as G1, 9 as G2, and 7 as G3.
2. By examining the intrinsic parameters of the clusters, we identify several clear trends. For example, G1 clusters typically exhibit lower total mass, lower escape velocity ($v_{\rm esc}$), lower concentration ($c$), and higher degree of tidal filling ($r_{\rm h,m}/r_{\rm t}^{\rm dyn}$), with statistically significant differences across morphological types. These patterns highlight the role of the gravitational potential well in suppressing stellar escape, while also emphasizing the internal fragility that makes clusters more prone to tidal disruption, even though a certain degree of overlap remains in both one- and two-dimensional parameter spaces.
@@
5. We discussed several observational and methodological limitations that may affect the identification of tidal structures, including data depth, smoothing scales, and the choice of tidal boundary. In this study, we mitigated these issues by adopting a homogeneous dataset from the DESI Legacy Imaging Surveys, applying a uniform matched-filtering and smoothing scheme, cross-validating our results against those from the literature, and explicitly considering both the King tidal radius and the dynamical tidal radius from authoritative globular cluster catalogues in our morphological classification. These measures effectively reduce systematic biases and enhance the reliability of tidal structure identification.
Overall, our study supports the view that the formation of extra-tidal structures is not governed by a single factor but arises from the complex interplay between a cluster’s internal properties and the external Galactic environment. Internal parameters, particularly those linked to the depth of the gravitational potential well, determine a cluster’s fundamental susceptibility to stellar stripping, while the external environment largely regulates the efficiency of this process. Together, these coupled effects shape both the likelihood of developing extra-tidal features and the morphological diversity observed among Galactic globular clusters.

### v2

@@ 4.3 Influence of orbital parameters on tidal feature formation
To further investigate the orbital dynamical mechanisms that may drive the formation of tidal structures in GCs, we analyzed a set of orbital parameters that are commonly linked to tidal interactions with the Galactic potential. Specifically, we considered the pericentric distance ($R_{{\rm Peri}}$), apocentric distance ($R_{{\rm Apo}}$), orbital semi-major axis ($a$), orbital inclination angle ($i$), orbital eccentricity ($e$), and the radial angle in action-angle coordinates ($\theta_{\rm R}$). These quantities collectively describe the cluster’s trajectory through the Galaxy and its exposure to varying tidal forces over time. All basic data are taken from the globular cluster database888people.smp.uq.edu.au/HolgerBaumgardt/globular/orbits.html. The semi-major axis, eccentricity, and inclination are computed following the formulae in Piatti (2019). The radial angle ($\theta_{\rm R}$) in action-angle coordinates is calculated using the Galpy package, adopting the MWPotential2014 Galactic potential (Bovy 2015). This parameter not only characterizes the current position of the cluster along its orbit but also indicates its direction of motion between the pericenter and apocenter. Specifically, when $0^{\circ}<\theta_{\rm R}<180^{\circ}$, the cluster is moving outward from pericenter toward apocenter, whereas when $180^{\circ}<\theta_{\rm R}<360^{\circ}$, it is moving inward from apocenter back toward pericenter.
@@ 4.3.1 Possible group-sensitive parameters
In the two-dimensional orbital parameter space, several notable correlations emerge. Clusters with smaller pericentric distances ($R_{\rm Peri}$) generally exhibit higher mass-loss fractions, consistent with the expectation that stronger tidal forces near the Galactic center accelerate mass depletion. This relation resembles a rough power-law trend, although some G3 clusters with small $R_{\rm Peri}$ show relatively low mass-loss fractions, suggesting that additional mechanisms may mitigate their long-term stripping. Clear geometric correlations are also evident among $R_{\rm Peri}$, $R_{\rm Apo}$, and $a$, but these simply reflect basic orbital mechanics. Clusters on highly eccentric ($e\gtrsim 0.7$) and strongly inclined ($i\gtrsim 70^{\circ}$) orbits are more frequently found among G1, and such orbital configurations are typically associated with enhanced mass loss, even though some overlap exists among different morphological groups. Webb et al. (2014) argued that repeated passages of GCs through the Galactic disk can amplify tidal disruption and increase mass loss. Building on this, Piatti (2019) proposed that clusters with higher eccentricities and inclinations are more likely to undergo repeated disk crossings, thereby increasing their susceptibility to tidal depletion. Piatti and Carballo-Bello (2020) further showed that clusters on highly eccentric ($e\gtrsim 0.8$) and strongly inclined ($\approx\pm 70^{\circ}$) orbits tend to lose more mass than those on nearly circular, low-inclination orbits. Our results are broadly consistent with this framework, while also indicating that not all clusters on such orbits necessarily display prominent tidal features.
@@
Moreover, Balbinot and Gieles (2018) suggested that GCs exhibiting tidal tails are typically in a near-dissolution phase and are more likely to be found near apocenter. In our sample, a considerable fraction of G1 clusters are located in the outward-moving orbital phase between pericenter and apocenter (with $\theta_{\rm R}$ between 0 and 180), consistent with the expectation that enhanced tidal stripping occurs during pericentric passages. However, we do not find clear evidence that these clusters are preferentially concentrated near apocenter. Although $\theta_{\rm R}$ and the mass-loss fraction show no strong correlation, G1 clusters nonetheless tend to populate this outward-moving phase, suggesting that the tidal stripping experienced during recent pericentric passages may contribute to the formation of their observed extratidal structures.
@@
Taken together, our results indicate that certain orbital characteristics, such as smaller $R_{\rm Peri}$, higher $e$, and specific $\theta_{\rm R}$, may be associated with enhanced tidal mass loss, and that these three parameters are group-sensitive. These findings highlight that, beyond intrinsic properties, the orbital configuration of GCs also plays a key role in determining both the extent of their mass loss and the detectability of their tidal structures.
4.3.2 Tidal feature orientations versus orbital direction
@@
Meanwhile, to investigate whether the orientation of tidal structures is related to the orbital direction of the clusters, we overlaid the orbital directions on the matched-filter S/N maps shown in Fig. 2. The sky-blue arrows indicate the short-period orbital directions derived from orbital integration, while the orange arrows mark the directions toward the Galactic center. The orbital integration was performed using the Galpy package with the MWPotential2014 gravitational model (Bovy 2015). This directional information provides additional context for interpreting whether the detected features are shaped primarily by orbital motion or Galactic tides.
@@
Our analysis reveals three distinct alignment patterns of tidal features among the clusters in our sample. In some systems (e.g., Whiting 1, NGC 5466), the extended structures broadly align with the orbital path, consistent with theoretical expectations that tidal stripping occurs preferentially along the orbit, producing leading and trailing tails (e.g., Montuori et al. 2007; Küpper et al. 2010). In others (e.g., NGC 1261, NGC 7492), the extensions are partially oriented toward the Galactic center, suggesting that the external tidal field has played a key role, possibly through recent disk-crossing events or similar perturbations (e.g., Combes et al. 1999; Balbinot et al. 2011). A third group of clusters (e.g., NGC 1851, NGC 6981) displays more symmetric, envelope-like morphologies with no clear alignment. In some cases, certain systems show visible tidal tails that are not aligned with the orbital direction (e.g., NGC 1904, NGC 4147), which may be related to recent pericentric passages (Awad et al. 2025).
@@
Figure 4: Corner plot of orbital parameters. The parameters included are pericenter distance ($R_{\mathrm{Peri}}$), apocenter distance ($R_{\mathrm{Apo}}$), orbital semi-major axis ($a$), inclination angle(ii), orbital eccentricity($e$), and radial angle in action-angle coordinates ($\theta_{\rm R}$). The mass-loss fraction ($M_{\rm dis}/M_{\rm init}$) is also included to test its potential coupling with orbital properties. The color-coding of morphological classifications, the use of symbols, and the presentation of diagonal/off-diagonal panels are the same as in Fig. 3.
@@ 4.4 Environmental independence of tidal morphologies
Progenitor: To assess whether accretion origin plays a role in shaping the outer structures of GCs, we examined their distribution in the energy-angular momentum ($E$-$L_{z}$) plane (Fig. 6), a diagnostic that has been widely applied to trace the assembly history of the Galactic halo and to connect clusters with past merger events (see, e.g., Massari et al. 2019; Naidu et al. 2020; Malhan et al. 2022). The associations between clusters and their progenitors adopted in this study are based on Massari et al. (2019, 2025 update) and Forbes (2020).
@@
Taken together, our analysis indicates that clusters with tidal features are distributed across a wide range of pericentric distances, orbital periods, pericentric passage times, and disk-crossing times, showing no clear preference for regions experiencing stronger or more frequent tidal forces. Likewise, accretion history provides no definitive evidence of driving the formation of extended tidal structures.
Figure 5: Orbital parameter relationships for clusters in different morphological groups. (a) The distribution of pericentric distance ($R_{\rm Peri}$) versus radial orbital period ($P_{\rm R}$). (b) The time since the last pericentric passage ($T_{\rm Peri}$) versus the time since the most recent disk crossing ($T_{\rm Disk}$). Clusters are color-coded by morphological group: G1 in red, G2 in blue, and G3 in black. Pal 12 is highlighted with an open marker.
Figure 6: Distribution of GCs in the $L_{z}$-$E$ space. Symbols indicate possible progenitor origins: circles represent clusters likely associated with the Gaia-Sausage-Enceladus (GSE) merger, triangles denote clusters linked to the Sagittarius (Sag) dwarf galaxy, squares indicate possible origin in the Sequoia (Seq) event, diamonds denote clusters associated with the Helmi stream (H99), and crosses mark clusters that may belong to either Gaia-Sausage-Enceladus or the Helmi stream (GSE/H99). Clusters are color-coded according to morphological classification: red for G1, blue for G2, and black for G3. Pal 12 is highlighted with open marker.
@@ 4.5 Limitations in the detection of tidal structures
Despite recent advances in wide-field surveys and data processing techniques, the identification of tidal structures around Galactic GCs remains subject to several observational and methodological limitations. These factors may introduce biases in the detection and interpretation of extended features, potentially leading to an underestimation of their prevalence, morphology, or physical origin.
@@
The data depth is crucial for assessing the completeness of tidal structures. In this work, we make use of imaging data from the DESI Legacy Survey, which reaches down to $r$ or $g$ $\sim$23-24 mag. For the GCs in our sample, this depth corresponds to 2-3 magnitudes below the main-sequence turnoff, thereby encompassing the majority of stellar populations associated with the clusters. Importantly, the Legacy Survey represents the deepest publicly available wide-field dataset that contains the largest number of GCs. Thus, the data employed here greatly reduces the biases arising from limited photometric depth, allowing for a more complete characterization of tidal features.
@@
Methods such as matched filtering (Rockosi et al. 2002) require careful choices of the spatial binning scale and the size of the smoothing kernel applied to density maps. Excessively large bins or overly broad kernels can make thin tidal tails difficult to discern, whereas overly small bins and kernels may introduce excessive noise and generate spurious structures. Therefore, by testing a range of spatial binning and smoothing scales and cross-checking our results against independent studies (see Appendix A), we minimize the risk of misclassification arising from parameter choices and ensure the reliability of our morphological classification.
@@
The tidal boundary of a globular cluster is often approximated by the King model tidal radius (King 1962), derived from fitting the surface density profile under the assumption of dynamical equilibrium. However, this estimate does not take into account orbital effects or the time-varying nature of the Galactic potential. Dynamical alternatives (such as the Jacobi radius and, in this work, the dynamical tidal radius at the current orbital phase) depend on the cluster’s mass, orbital phase, and the adopted Galactic potential model. Using different $r_{\rm t}$ values can change the classification of outer stars as ”bound” or ”extratidal”, thereby affecting whether a cluster is deemed to possess extended structures.
To mitigate these effects, we adopted both the widely used $r_{\rm t}^{\rm King}$ (Harris 1996, 2010 edition) from the literature and $r_{\rm t}^{\rm dyn}$ (from the globular cluster databases referenced in this work) calculated from orbital parameters. Both sets of values are taken from well-established and authoritative cluster catalogs, ensuring consistency with previous studies. In our morphological classification, we explicitly considered both definitions, thereby reducing the limitations associated with relying on a single tidal radius and enhancing the reliability of extratidal structure identification.
@@
In summary, although matched filtering and wide-field imaging have improved our ability to identify extended structures in GCs, the detection remains vulnerable to various systematic effects. In this study, we minimized these effects by employing a homogeneous dataset, adopting uniform filtering parameters and processing procedures, and cross-validating our results with previous imaging studies. These measures reduce potential systematic biases and enhance the reliability of the identified tidal features and their morphological classifications.
@@ 5 Summary
In this study, we conducted a systematic investigation of the extratidal structures associated with Galactic GCs of likely extragalactic origin, aiming to explore their formation mechanisms and the key physical factors driving their morphological diversity. Using a homogeneous dataset from the DESI Legacy Imaging Surveys and a uniform matched-filtering procedure, we identified and morphologically classified the tidal features of our cluster sample. We then explored the connections between extratidal structures and clusters’ intrinsic properties, orbital dynamical parameters, and Galactic environment, followed by a discussion of several observational and methodological limitations affecting their detection and classification. Within this framework, we arrived at the following results:
1. Our analysis shows that a substantial fraction of clusters exhibit extended tidal structures, including prominent tidal tails (G1) and diffuse stellar envelopes (G2), which in many cases extend well beyond both the King tidal radius and the dynamical tidal radius. In contrast, some clusters remain spatially confined within their nominal tidal boundaries and show no evidence of extratidal features (G3). In total, 12 clusters are classified as G1, 9 as G2, and 7 as G3.
2. By examining the intrinsic parameters of the clusters, we identified several clear trends. For example, G1 clusters typically exhibit lower total mass, lower escape velocity ($v_{\rm esc}$), lower concentration ($c$), and higher degree of tidal filling ($r_{\rm h,m}/r_{\rm t}^{\rm dyn}$), with statistically significant differences across morphological types. These patterns highlight the role of the gravitational potential well in suppressing stellar escape, while also emphasizing the internal fragility that makes clusters more prone to tidal disruption, even though a certain degree of overlap remains in both one- and two-dimensional parameter spaces.
@@
5. We discussed several observational and methodological limitations that may affect the identification of tidal structures, including data depth, smoothing scales, and the choice of tidal boundary. In this study, we mitigated these issues by adopting a homogeneous dataset from the DESI Legacy Imaging Surveys, applying a uniform matched-filtering and smoothing scheme, cross-validating our results against those from the literature, and explicitly considering both the King tidal radius and the dynamical tidal radius from authoritative globular cluster catalogs in our morphological classification. These measures effectively reduce systematic biases and enhance the reliability of tidal structure identification.
Overall, our study supports the view that the formation of extratidal structures is not governed by a single factor but arises from the complex interplay between a cluster’s internal properties and the external Galactic environment. Internal parameters, particularly those linked to the depth of the gravitational potential well, determine a cluster’s fundamental susceptibility to stellar stripping, while the external environment largely regulates the efficiency of this process. Together, these coupled effects shape both the likelihood of developing extratidal features and the morphological diversity observed among Galactic GCs.